OpenStudy (ittom14):

During a baseball game, a batter hits a high pop-up. The ball remains in the air for 6.0 seconds. How high does it rise?

OpenStudy (jiteshmeghwal9):

Let \(s_y\) be the displacement of ball along y-direction which is 0 as the ball came back to its initial position. And \(u_y\) be the initial velocity of ball. The ball is moving under constant force of gravity so we may apply equations of motion here.

OpenStudy (jiteshmeghwal9):

\[s_y=u_yt-\frac{1}{2} gt^2\] i assumed downward direction negative and upward direction to be positive.

OpenStudy (jiteshmeghwal9):

\[0=6u_y-\frac{1}{2} g(6)^2\]\[u_y=30 \]

OpenStudy (jiteshmeghwal9):

Now we found the initial velocity of ball . The ball reached its maximum height when the velocity of ball becomes zero. \[v_y^2=u_y^2-2gh_{max}\] \(v_y=0\) so \[0=(30)^2-2gh_{max}\] find \(h_{max}\) which is maximum height achieved.